### v1

@@ B.1 Lemma
Next we give a bound on $N_{\varepsilon}$. It is a classical result (see for instance [24]) that the covering number of the unit ball $B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})):=\{u\in C^{2}(\mathbb{I}^{d_{N_{j}}}):\left\|u\right\|_{C^{2}}\leq 1\}$ is bounded by
$\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})),L^{\infty})\leq C_{d_{N_{j}}}\varepsilon^{-d_{N_{j}}/2}.$
Here $\mathcal{U}_{\theta}$ is the product space of $C^{2}(\mathbb{I}^{d_{N_{j}}})$ equipped with the $\ell_{\infty}$-norm, so that
$\log N_{\varepsilon}\leq\sum_{j\in V}\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})),L^{\infty})\leq bC_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}.$
Recall here $d_{\rm loc}=\max_{j}d_{N_{j}}$. Notice $f_{\theta,j}$ is uniformly bounded by
$|f_{\theta,j}|\leq 2(s+1)R\sum_{k\in N_{j}}\left\|u_{\theta,k}\right\|_{C^{2}}\leq 2(s+1)^{2}R.$
@@
$\mathbb{P}\left[|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq\mathrm{e}^{-Nt^{2}/2(s+1)^{4}R^{2}},\quad\forall t>0.$
@@
$\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-Nt^{2}/2(s+1)^{4}R^{2}}.$
Take $t=(s+1)^{2}R\varepsilon$, then
@@
Combine the above estimate, we obtain
$\mathbb{P}\left[J_{\rm loc}(\hat{\theta}_{*})\geq 6(s+1)^{2}R^{2}\varepsilon\right]\leq b^{2}C_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}\mathrm{e}^{-N\varepsilon^{2}/2}.$
In order to make the right hand side less than some $\delta>0$, one can take
$\varepsilon^{2}=\frac{C}{N}\left(\log\delta^{-1}+\log b\right),$
where $C$ is some constant depending only on $d_{\rm loc}$. So that with probability at least $1-\delta$,
$J_{\rm loc}(\hat{\theta}_{*})\leq C^{\prime}N^{-1/2}\sqrt{\log\delta^{-1}+\log b}.$
The conclusion follows by plugging in the bound (C.1) on $\max_{j}W_{1}(p_{j},\hat{p}_{j})$.
Appendix D Proof for 5.1
The proof is based on the following approximation result adapted from Theorem 3 in [28]:
Proposition D.1.
For any $u\in W^{2,\infty}(\mathbb{I}^{d_{0}})$ and $\left\|u\right\|_{W^{2,\infty}}\leq 1$, there exists a neural network $u_{\theta}\in\Phi_{d_{0}}(L,W)$ where $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ such that
$\left\|u-u_{\theta}\right\|_{W^{1,\infty}(\mathbb{I}^{d_{0}})}\leq C_{3}L_{0}^{-2/d_{0}}W_{0}^{-2/d_{0}}.$
Here $C_{1},C_{2},C_{3}$ are constants depending only on $d_{0}$.
By 2.7, for $p_{\theta}\in\mathscr{P}_{\rm NN}$, it holds that (denote $\delta=\frac{sr!\kappa^{r}}{m}$)
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq\delta\cdot\max_{k\in V}\left\|\nabla_{k}\log p-\nabla_{k}\log p_{\theta}\right\|_{L^{1}(p)}.$
By the sprase graphical structure,
$\nabla_{k}\log p-\nabla_{k}\log p_{\theta}=\sum_{v\in N_{k}}\left(\nabla_{k}u_{j}(x_{N_{j}})-\nabla_{k}u_{\theta,j}(x_{N_{j}})\right).$
Therefore, since $G$ is $s$-sparse,
$\left\|\nabla_{k}\log p-\nabla_{k}\log p_{\theta}\right\|_{L^{1}(p)}\leq$ $\sum_{v\in N_{k}}\left\|\nabla_{k}u_{j}-\nabla_{k}u_{\theta,j}\right\|_{L^{1}(p)}$
$\leq(s+1)\max_{j\in V}$ $\left\|\nabla_{k}u_{j}-\nabla_{k}u_{\theta,j}\right\|_{L^{\infty}}\leq(s+1)\max_{j\in V}\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}.$
We obtain
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq\delta(s+1)\cdot\max_{j\in V}\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}.$
By D.1, we can take $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ in $\mathscr{P}_{\rm NN}$, so that for any $j\in V$, there exists $u_{\theta,j}\in\Phi_{d_{N_{j}}}(L,W)$ s.t. (notice we take $\left\|u_{j}\right\|_{W^{2,\infty}}\leq R$)
$\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}\leq C_{3}RL_{0}^{-2/d_{N_{j}}}W_{0}^{-2/d_{N_{j}}}\leq C_{3}RL_{0}^{-2/d_{\rm loc}}W_{0}^{-2/d_{\rm loc}}.$
Here $d_{\rm loc}=\max_{j\in V}d_{N_{j}}$, and the constant $C_{1},C_{2},C_{3}$ depend only on $d_{\rm loc}$. This completes the proof.

### v2

@@ B.1 Lemma
Next we give a bound on $N_{\varepsilon}$. It is a classical result (see for instance [43]) that the covering number of the unit ball $B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})):=\{u\in C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}}):\left\|u\right\|_{C^{2}}\leq 1\}$ is bounded by
$\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})),L^{\infty})\leq C_{d_{\mathcal{N}_{j}}}\varepsilon^{-d_{\mathcal{N}_{j}}/2}.$
Here $\mathcal{U}_{\theta}$ is the product space of $C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})$ equipped with the $\ell_{\infty}$-norm, so that
$\log N_{\varepsilon}\leq\sum_{j\in V}\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})),L^{\infty})\leq bC_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}.$ (C.3)
Recall here $d_{\rm loc}=\max_{j}d_{\mathcal{N}_{j}}$. Notice $f_{\theta,j}$ is uniformly bounded by
$|f_{\theta,j}|\leq 2(S+1)R\sum_{k\in\mathcal{N}_{j}}\left\|u_{\theta,k}\right\|_{C^{2}}\leq 2(S+1)^{2}R^{2}.$
@@
$\mathbb{P}\left[|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq\mathrm{e}^{-Nt^{2}/2(S+1)^{4}R^{4}},\quad\forall t>0.$
@@
$\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-Nt^{2}/2(S+1)^{4}R^{4}}.$
Take $t=(S+1)^{2}R^{2}\varepsilon$, then
@@
Recall (C.2) and (C.3), we obtain
$\mathbb{P}\left[J_{\rm loc}(\hat{\theta}_{*})\geq 6(S+1)^{2}R^{2}\varepsilon\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-N\varepsilon^{2}/2}\leq b^{2}C_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}\mathrm{e}^{-N\varepsilon^{2}/2}.$
In order to make the right hand side less than $\lambda\in(0,1)$, one can take
$\varepsilon^{2}=\frac{C}{N}\left(\log\lambda^{-1}+\log b\right),$
where $C$ is some constant depending only on $d_{\rm loc}$. So that with probability at least $1-\lambda$,
$J_{\rm loc}(\hat{\theta}_{*})\leq 6(S+1)^{2}R^{2}\varepsilon\leq C^{\prime}N^{-1/2}\sqrt{\log\lambda^{-1}+\log b}.$
The conclusion follows by plugging in the bound (C.1) on $\max_{j}\mathsf{W}_{1}(p_{j},\hat{p}_{j})$. ∎